# Data Diet: Can Trimming PET/CT Datasets Enhance Lesion Segmentation?

### v2

@@ 2.4 Proposed Method
According to our previous analysis, we find, that an ideal, data-centric solution should focus on reducing the predicted false positive volumes of the model for PSMA-PETs. With our intent of selecting samples to exclude from the training set, we reduce the search space for such samples to the PSMA-PETs and keep all FDG-PET images as FDG tracer images scored already quite high in our measurements. Regarding the PSMA-PETs we exclude images for which the baseline model produces a very low loss, which are potential candidates that tilt the model towards overconfidence, as the model overfits on them. To this end, we use the measured loss from the previous step and order the PSMA-images by their loss magnitude in ascending order, then we prune the $n^{\text{th}}$ percentile from the training dataset before retraining from scratch, where $n$ is a hyperparameter that we ablate. When analyzing the PSMA images that produce very low loss values ($n\leq 5$), we find not a single image of them contains a healthy patient. This approach that we call data diet poses two advantages: (1) It improves the sick-to-healthy ratio of the training dataset and (2) hard examples – where learning happens – remain in the training dataset which should lead to a better calibrated model and thus reduce the false positives.
@@ 3 Results
We find that our established intuition holds quite well with the false negative volume generally decreasing with increasing data exclusion. The false positive volume tends to increase as well, which is reasonable, given the model becomes more cautious when predicting metastasis. Taking a look at the Dice score, which is driven by both, false positives as well as false negatives, we find the best results when excluding $3\%$ of the easiest PSMA images. We give the full list of excluded images for the $5\%$ and the $3\%$ scenario in Appendix 0.A. To further evaluate whether the proposed approach reduced the number of false positive predictions in the PSMA-Tracer, we generated a QQ-plot of the log-percentiles from the PSMA False-Positive Distribution, comparing the training dataset before and after a 3% data diet. The results are presented in fig. 3. The plot reveals that most quantiles of the distribution shifted below the dashed reference line, which indicates that the pre- and post-diet distributions are equal. This shift demonstrates a reduction in false positive values, supporting the effectiveness of the data diet approach. Interestingly, the data diet did not alter the False-Negative Volume distribution, as shown in a similar log QQ-plot in Appendix 0.C.
@@ 4 Discussion and Limitation
While the overall hypothesis seems to match the observed results it has to be noted that due to the rather small size of the preliminary test set consisting only of a handful of patients, further experiments would be required to strengthen the empirical evidence. Further, we rely on the assumption that modifying the PSMA-data does not influence the performance of the FDG-PET images which would require additional experiments to be validated or falsified.

### v4

@@ 2.4 Proposed Method
According to our previous analysis, we find, that an ideal, data-centric solution should focus on reducing the predicted false positive volumes of the model for PSMA-PETs. With our intent of selecting samples to exclude from the training set, we reduce the search space for such samples to the PSMA-PETs and keep all FDG-PET images as FDG tracer images scored already quite high in our measurements. Regarding the PSMA-PETs we exclude images for which the baseline model produces a very low loss, which are potential candidates that tilt the model towards overconfidence, as the model overfits on them. To this end, we use the measured loss from the previous step and order the PSMA-images by their loss magnitude in ascending order, then we prune the $n^{\text{th}}$ percentile from the training dataset before retraining from scratch, where $n$ is a hyperparameter that we ablate. When analyzing the PSMA images that produce very low loss values ($n\leq 5$), we find not a single image of them contains a healthy patient. This approach that we call data diet poses two advantages: (1) It improves the sick-to-healthy ratio of the training dataset and (2) hard examples – where learning happens – remain in the training dataset which should lead to a better-calibrated model and thus reduce the false positives.
@@ 3 Results
Taking a look at the Dice score, which is driven by both, false positives as well as false negatives, we find the best results when excluding $3\%$ of the easiest PSMA images, we however find that due to the limited size of the preliminary testset (only 5 images) we can neither confirm nor contradict our proposed strategy. We give the full list of excluded images for the $5\%$ and the $3\%$ scenario in Appendix 0.A. To further evaluate whether the proposed approach reduced the number of false positive predictions in the PSMA-Tracer, we generated a QQ-plot of the log-percentiles from the PSMA False-Positive Distribution, comparing the training dataset before and after a 3% data diet. The results are presented in fig. 3. The plot reveals that most quantiles of the distribution shifted below the dashed reference line, which indicates that the pre- and post-diet distributions are equal. This shift demonstrates a reduction in false positive values, supporting the effectiveness of the data diet approach. Interestingly, the data diet did not alter the False-Negative Volume distribution, as shown in a similar log QQ-plot in Appendix 0.C.
@@ 4 Discussion and Limitation
While the overall hypothesis seems to match the observed results it has to be noted that due to the rather small size of the preliminary test set consisting only of a handful of patients, further experiments would be required to confirm the empirical evidence. Further, we rely on the assumption that modifying the PSMA-data does not influence the performance of the FDG-PET images which would require additional experiments to be validated or falsified.